OpenStudy (anonymous):

If sin(θ) = x/8 for − π/2 < θ < π/2 , find an expression for cos(2θ) in terms of x.

OpenStudy (anonymous):

is x\8 dividing or a fraction?

OpenStudy (anonymous):

fraction

OpenStudy (dumbcow):

\[\cos 2\theta = \cos^{2} \theta - \sin^{2} \theta = 1-2\sin^{2} \theta\] plug in "x/8" for sine